OpenStudy (anonymous):

how is 50/100 get to 1/2

OpenStudy (anonymous):

it is basically the same thing 50/100 =1/2 if you divide the numerator by 50(50/50=1) you would get 1 what you do to the top you do to the bottom so divide 100/50=2 so 50/100 = 1/2.